# Large Language Models for Anomaly and Out-of-Distribution Detection: A Survey

### v1

@@ Abstract
Detecting anomalies or out-of-distribution (OOD) samples is critical for maintaining the reliability and trustworthiness of machine learning systems. Recently, Large Language Models (LLMs) have demonstrated their effectiveness not only in natural language processing but also in broader applications due to their advanced comprehension and generative capabilities. The integration of LLMs into anomaly and OOD detection marks a significant shift from the traditional paradigm in the field. This survey focuses on the problem of anomaly and OOD detection under the context of LLMs. We propose a new taxonomy to categorize existing approaches into three classes based on the role played by LLMs. Following our proposed taxonomy, we further discuss the related work under each of the categories and finally discuss potential challenges and directions for future research in this field. We also provide an up-to-date reading list  [^1] https://github.com/rux001/Awesome-LLM-Anomaly-OOD-Detection of relevant papers.
@@ 1 Introduction
Figure 1: A simple illustration of leveraging LLMs for images anomaly and OOD detection.
Large Language Models (LLMs), such as GPT-4 Achiam et al. (2023) and LLaMA Touvron et al. (2023), have demonstrated remarkable capabilities in language comprehension and summarization. To further harness the potential of LLMs beyond text data, there is also a growing interest in extending them to multi-modal tasks such as vision-language understanding and generation Wang et al. (2024), evolving them into Multimodal LLMs (MLLMs) Yin et al. (2023). Given the zero- and few-shot reasoning capabilities of LLMs and MLLMs, researchers try to apply these models to anomaly and out-of-distribution (OOD) detection, as illustrated in Figure 1, yielding promising detection results.
@@
In this survey, we propose a novel taxonomy that focuses on how LLMs can profoundly impact anomaly and OOD detection in three fundamental ways, as illustrated in Figure 2: ❶ LLMs for Augmentation (§3): LLMs are not used directly for detection, but their emergent abilities, advanced semantic understanding, and vast knowledge augment the detection process; ❷ LLMs for Detection (§4): LLMs are employed as a detector to identify anomalies and OOD instances; and ❸ LLMs for Explanation (§5): LLMs provide insightful explanatory analyses of detection results, aiding in further planning and problem-solving in real-world scenarios. At the end (§6), we also outline the challenges and future research directions, in order to provide a better understanding of anomaly and OOD detection in the era of LLMs and shed light on the following research.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=6.4em,font=,, where level=2text width=4.9em,font=,, where level=3text width=4.9em,font=,, where level=4text width=6.1em,font=,, [ LLMs for Anomaly and OOD Detection, ver [ LLMs for Augmentation [ Text Embedding-based,text width = 7.6em [ LogGPT Qi et al. (2023), LogFit Almodovar et al. (2024),  Liu et al. (2024a),  Zhang et al. (2024a), leaf, text width=30.35em ] ] [ Psuedo Label-based,text width = 7.6em [ EOE Cao et al. (2024), PCC Huang et al. (2024b), TOE Park et al. (2023), CoNAL Xu et al. (2023), leaf, text width=30.35em ] ] [ Textual Description-based,text width = 7.6em [ TagFog  Chen et al. (2024), ALFA Zhu et al. (2024),  Dai et al. (2023), leaf, text width=30.35em ] ] ] [ LLMs for Detection [ Prompting,text width=3em, middle [ w/o Tuning,text width=3.2em,middle [ SIGLLM Alnegheimish et al. (2024), LLMAD Liu et al. (2024c), LogPrompt Liu et al. (2024d), LAVAD
 Zanella et al. (2024), LLM-Monitor Elhafsi et al. (2023), GPT-4V-AD  Zhang et al. (2023), Cao et al. (2023), leaf, text width=30.5em ] ] [ w/ Tuning,text width=3.2em,middle [ Tabular Li et al. (2024a), Myriad Li et al. (2023b), AnomalyGPT Zhang et al. (2023), leaf, text width=30.5em ] ] ] [ Contrasting,text width=3em,middle [ w/o Tuning,text width=3.2em,middle [ Fort et al. (2021), ZOC Esmaeilpour et al. (2022), NegLabel Jiang et al. (2024), CLIPScope Fu et al. (2024),
WinCLIP  Jeong et al. (2023), AnoCLIP Deng et al. (2023), CLIP-AD Chen et al. (2023), MCM
 Ming et al. (2022),  Miyai et al. (2023), SETAR Li et al. (2024c), leaf, text width=30.5em ] ] [ w/ Tuning,text width=3.2em,middle [ LoCoOp Miyai et al. (2024c) AnomalyCLIP Zhou et al. (2024), InCTRL Zhu and Pang (2024), MVFA
 Huang et al. (2024a), ID-like Bai et al. (2024), NegPrompt Li et al. (2024b), CLIPN Wang et al. (2023),
LSN Nie et al. (2024), MCM-PEFT Ming and Li (2024), leaf, text width=30.5em ] ] ] ] [ LLMs for Explanation [ Holmes-VAD Zhang et al. (2024b), AnomalyRuler Yang et al. (2024b), VAD-LLaMA  Lv and Sun (2024), AESOP Sinha et al. (2024), leaf, text width=39.5em ] ] ]
@@ 2.1 Problem Definition
3 LLMs for Augmentation
In this section, we review methods that leverage LLM as a data augmenter, producing meaningful augmented knowledge that enhances the detection of anomalies or OOD samples. Such augmented information includes text embedding, pseudo labels, and textual descriptions derived from LLMs. Therefore, these approaches can be categorized into three types as shown in Figure 3.
3.1 Text Embedding-based Augmentation
LLMs are powerful feature extractor which can derive meaningful and effective embedding used for further detection tasks. For instance, in log data, Hadadi et al. (2024) and Qi et al. (2023) fine-tune pre-trained GPT models in a supervised manner and use the extracted semantic embeddings as an important component for future anomaly detection.
For OOD detection in text data, a standard pipeline involves using encoder-only LLMs to generate sentence representations, which are then used to derive OOD confidence scores. Typically, these models are fine-tuned on ID data, and OOD detectors are applied to the sentence representations they produce Liu et al. (2024a). Recently, there has been a shift toward leveraging larger language models with decoder architectures, which offer enhanced capabilities in extracting and refining textual representations. Liu et al. (2024a) explore the use of decoder-only LLMs, such as LLaMa, incorporating fine-tuning techniques like LoRA to minimize additional parameter usage. Their findings demonstrate that fine-tuned LLMs, when combined with customized OOD scoring functions, can significantly improve OOD detection performance. A key advantage of recent LLMs with decoder architecture is their autoregressive ability, which allows for more effective handling of sequential data. Building on this, Zhang et al. (2024a) propose using the likelihood ratio between a pre-trained LLM and its fine-tuned variant as a criterion for OOD detection, effectively leveraging the deep, contextual knowledge embedded within LLMs for text data.
3.2 Pseudo Label-based Augmentation
The emergent abilities of LLMs offer a promising approach for generating high-quality synthetic datasets that, in some cases, can surpass those curated by humans Ding et al. (2024). A significant challenge in using LLMs for OOD detection is the lack of OOD labels, which often hampers model performance. Traditional methods rely on extensive human effort and auxiliary datasets, but LLMs can overcome this by generating high-quality pseudo-OOD labels through suitable prompts. These pseudo labels can then be used as text prompts for contrasting-based OOD detection methods, augment existing ID data and enhance the distinction between ID and OOD samples during detection.
Figure 3: The illustration of three approaches in (§3): (a) Text Embedding-based Augmentation; (b) Pseudo Label-based Augmentation, and (c) Textual Description-based Augmentation.
EOE Cao et al. (2024) and PCC Huang et al. (2024b) prompt LLMs to generate potential OOD class labels which are visually similar to known ID classes. Then, they define a new score function with penalty on these generated pseudo labels during inference stage, greatly outperforming methods with only known ID labels. Following the similar idea, TOE Park et al. (2023) further evaluates generating pseudo OOD labels for OOD detection at three verbosity levels: word-level, description-level, and caption-level, using BERT, GPT-3 and BLIP-2 respectively. Results indicate that using caption-level pseudo OOD labels outperform other two approaches since BLIP-2 can leverage both semantic and visual underdtanding. For text data, CoNAL Xu et al. (2023) prompts LLMs to extend closed-set labels with novel ones and generates new examples based on these labels, forming a comprehensive set of probable OOD samples. By utilizing a contrastive confidence loss for training, detection model achieves both high accuracy on the ID training set and lower relative confidence on the generated novel examples.
3.3 Textual Description-based Augmentation
In addition to generating pseudo lebels, other methods utilize LLMs to generate detailed textual descriptions about both known ID classes and potential unknown OOD samples. For example, TagFog Chen et al. (2024) uses the Jigsaw strategy to generate fake OOD samples and prompts ChatGPT to create detailed descriptions for each ID class, guiding the training of the image encoder of CLIP for OOD detection. When using LLMs for anomaly detection, it is crucial to make LLMs recognize the close correlation between normal images and their respective normal prompts, while identifying a more distant association with abnormal prompts. Therefore, detailed and nuanced descriptions of normal and anomalous stages of an object are necessary. ALFA Zhu et al. (2024) formulates prompts to query an LLM to describe normal and abnormal features for each class and then used these descriptions together as prompts for LLMs to better identify abnormal object. To avoid LLM hallucination issues, Dai et al. (2023) use LLMs to describe visual features for distinguishing categories in images and introduce a consistency-based uncertainty calibration method to estimate the confidence score of each generation.
@@ 4 LLMs for Detection
The primary objective of this section is to explore existing works that utilize LLMs to detect anomalies or OOD samples. Under this line of research, approaches can be categorized into two classes as illustrated in Figure 4: ❶ Prompting-based Detection methods, which involve directly prompting LLMs to generate language responses that include detection results; ❷ Contrasting-based Detection methods, which focus on multimodal scenarios, using MLLMs pre-trained with a contrastive objective as detectors.
@@ 4.1.1 Detection without LLM Tuning
Since some approaches do not require additional tuning, they mainly focus on employing various prompt engineering techniques Sahoo et al. (2024) to guide LLMs to produce better detection results. To design suitable prompts for anomaly or OOD detection, researchers have employed a combination of various prompt techniques, such as role-play prompting Wu et al. (2023), in-context learning Brown et al. (2020), and chain-of-thought (CoT) reasoning Wei et al. (2022), to create effective prompt templates. Studies such as SIGLLM Alnegheimish et al. (2024), LLMAD Liu et al. (2024c), and LogPrompt Liu et al. (2024d) focus on time series and log data. SIGLLM Alnegheimish et al. (2024) investigates two distinct pipelines for using LLMs in time series anomaly detection: one directly prompts an LLM with specific role-play instructions to identify anomalous elements in given data, and the other uses the LLM’s forecasting ability to detect anomalies by comparing original and forecasted signals, where discrepancies indicate anomalies. LLMAD Liu et al. (2024c) incorporates in-context learning examples retrieved from a constructed database and CoT prompts that inject domain knowledge of time series. LogPrompt Liu et al. (2024d) explores three prompting strategies for log data: self-prompt, CoT prompt, and in-context prompt, demonstrating that the prompt with CoT techniques outperforms other prompting strategies. The tailored CoT prompt for log data includes a specific task instruction, i.e. “classify the given log entries into normal and abnormal categories”, and step-by-step rules for considering given data as anomalies.
@@ 4.2.1 Detection without LLM Tuning
Despite the promise, existing CLIP-like models perform zero-shot classification in a closed-world setting. That is, it will match an input into a fixed set of categories, even if it is irrelevant Ming et al. (2022). To address this, one approach involves designing effective post-hoc score functions tailored for OOD detection that solely rely on ID class labels. Alternatively, some researchers incorporate anomaly or OOD class information into the text prompts, allowing the model to match OOD or abnormal images to paired prompts.
  • Without Anomaly/OOD Prompts. To address the challenges of OOD detection using only in-distribution (ID) class information while avoiding the matching of OOD inputs to irrelevant ID classes, one notable approach is the Maximum Concept Matching (MCM) framework proposed by Ming et al. (2022). This method is not limited to CLIP and can be generally applicable to other pre-trained models that promote multi-modal feature alignment. They view the textual embeddings of ID classes as a collection of concept prototypes and define the maximum concept matching (MCM) score based on the cosine similarity between the image feature and the textual feature. Following the idea of MCM, several subsequent works focus on improving OOD detection results by either adding a local MCM score or modifying weights in the original MCM framework, such as Miyai et al. (2023) and Li et al. (2024c).
@@
    For contrasting-based anomaly detection, WinCLIP Jeong et al. (2023) initially investigates a one-class design by using only the normal prompt “normal [o]” where [o] represents object-level label, i.e “bottle”, and defining an anomaly score as the similarity between vectors derived from the image encoder and normal prompts. However, this one-class design yields poorer results compared to a simple binary zero-shot framework, CLIP-AC Jeong et al. (2023), which adapts CLIP with two class prompts: “normal [o]” vs. “anomalous [o]”. This framework sets the foundational pipeline for future work in contrasting-based anomaly detection and has inspired many subsequent research.
    While using the default prompt has demonstrated promising performance, similar to the prompt engineering discussion around GPT-3 Brown et al. (2020), researchers have observed that performance can be significantly improved by customizing the prompt text. Models like WinCLIP Jeong et al. (2023) and AnoCLIP Deng et al. (2023) use a Prompt Ensemble technique to generate all combinations of pre-defined lists of state words per label and text templates. After generating all combinations of states and templates, they compute the average of text embeddings per label to represent the normal and anomalous classes. In practice, more descriptions in prompts do not always yield better performance. Therefore, CLIP-AD Chen et al. (2023) proposes Representative Vector Selection (RVS), from a distributional perspective for the design of the text prompt, broadening research opportunities beyond merely crafting adjectives.
@@ 4.2.2 Detection with LLM Tuning
  • LLM Adapter-Tuning. Adapter-tuning methods involve integrating additional components or layers into the model architecture to facilitate better alignment or localization Hu et al. (2023). CLIP was originally designed for classifying the semantics of objects in the scene, which does not align well with the sensory anomaly detection task where both normal and abnormal samples are often from the same class of object. To reconcile this, InCTRL Zhu and Pang (2024) includes a tunable adapter layer to further adapt the image representations for anomaly detection. To better adapt to medical image anomaly detection, MVFA Huang et al. (2024a) proposes a multi-level visual feature adaptation architecture to align CLIP’s features with the requirements of anomaly detection in medical contexts. This is achieved by integrating multiple residual adapters into the pre-trained visual encoder, guided by multi-level, pixel-wise visual-language feature alignment loss functions.
@@
    Instead of focusing on leveraging OOD information into the text encoder, some methods aim to perform prompt tuning to optimize word embeddings for ID labels and then use the MCM score as the detection criterion. MCM-PEFT Ming and Li (2024) demonstrates that simply applying prompt tuning for CLIP on few-shot ID datasets can significantly improve detection accuracy. However, a primary limitation of this approach is its exclusive reliance on the features of ID classes, leading to incorrect detection when input images share a high visual similarity with the class in the prompt. To address this, LoCoOp Miyai et al. (2024c) treats such ID-irrelevant nuisances as OOD and learns to push them away from the ID class text embeddings, preventing the model from producing high ID confidence scores for the OOD features. Additionally, Lafon et al. (2024) enhances detection capabilities by learning a diverse set of prompts utilizing both global and local visual representations. To better adapt to learning local features, AnomalyCLIP Zhou et al. (2024) aims to learn object-agnostic text prompts that capture generic normality and abnormality in images, allowing the model to focus on abnormal regions rather than object semantics.
5 LLMs for Explanation
Due to their remarkable capabilities in understanding and generating human-like text, LLMs have been explored for providing insightful explanations and analyses for anomaly and OOD detection results, thereby aiding in further planning, problem-solving and decision-making.

### v2

@@ Abstract
Detecting anomalies or out-of-distribution (OOD) samples is critical for maintaining the reliability and trustworthiness of machine learning systems. Recently, Large Language Models (LLMs) have demonstrated their effectiveness not only in natural language processing but also in broader applications due to their advanced comprehension and generative capabilities. The integration of LLMs into anomaly and OOD detection marks a significant shift from the traditional paradigm in the field. This survey focuses on the problem of anomaly and OOD detection under the context of LLMs. We propose a new taxonomy to categorize existing approaches into two classes based on the role played by LLMs. Following our proposed taxonomy, we further discuss the related work under each of the categories and finally discuss potential challenges and directions for future research in this field. We also provide an up-to-date reading list  [^1] https://github.com/rux001/Awesome-LLM-Anomaly-OOD-Detection of relevant papers.
@@ 1 Introduction
Figure 1: A simple illustration of leveraging LLMs for vision anomaly and OOD detection.
Large Language Models (LLMs), such as GPT-4 Achiam et al. (2023) and LLaMA Touvron et al. (2023), have demonstrated remarkable capabilities in language comprehension and generation. To further harness the potential of LLMs beyond text data, there is also a growing interest in extending them to multi-modal tasks such as vision-language understanding and generation Wang et al. (2024), evolving them into Multimodal LLMs (MLLMs) Yin et al. (2023). Given the zero- and few-shot reasoning capabilities of LLMs and MLLMs, researchers try to apply these models to anomaly and out-of-distribution (OOD) detection, as illustrated in Figure 1, yielding promising detection results.
@@
In this survey, we propose a novel taxonomy that focuses on how LLMs can profoundly impact anomaly and OOD detection in two fundamental ways, as illustrated in Figure 4: ❶ LLMs for Detection (§3): We provide a detailed review of existing methods that leverage LLMs as detectors for identifying anomalies and OOD instances, categorizing these methods into distinct pipelines; and ❷LLMs for Generation (§4): We also review methods that utilize LLMs’ emergent abilities, advanced semantic understanding, and vast knowledge to generate augmented data and explanations, thereby enhancing detection models with more robust and interpretable outputs. At the end (§5), we also outline the challenges and future research directions, in order to provide a better understanding of anomaly and OOD detection in the era of LLMs and shed light on the following research.
@@ 3 LLMs for Detection
The primary objective of this section is to explore existing works that utilize LLMs’ inherent knowledge to detect anomalies or OOD samples. Under this line of research, approaches can be categorized into two classes as illustrated in Figure 2: ❶ Prompting-based Detection methods, which involve directly prompting LLMs to generate language responses that include detection results; ❷ Contrasting-based Detection methods, which focus on multimodal scenarios, using MLLMs pre-trained with a contrastive objective as detectors.
@@ 3.1 Prompting-based Detection
Note that the prompting-based approach primarily addresses the anomaly detection task. OOD detection research has not yet widely adopted prompting to directly identify OOD samples.
@@ 3.1.1 Detection without LLM Tuning
Since some approaches do not require additional tuning, they mainly focus on employing various prompt engineering techniques Sahoo et al. (2024) to guide LLMs to produce better detection results. To design suitable prompts for anomaly detection, researchers have employed a combination of various prompt techniques, such as role-play prompting Wu et al. (2023), in-context learning Brown et al. (2020), and chain-of-thought (CoT) reasoning Wei et al. (2022), to create effective prompt templates. Studies such as SIGLLM Alnegheimish et al. (2024), LLMAD Liu et al. (2024c), and LogPrompt Liu et al. (2024d) focus on time series and log data. SIGLLM Alnegheimish et al. (2024) investigates two distinct pipelines for using LLMs in time series anomaly detection: one directly prompts an LLM with specific role-play instructions to identify anomalous elements in given data, and the other uses the LLM’s forecasting ability to detect anomalies by comparing original and forecasted signals, where discrepancies indicate anomalies. LLMAD Liu et al. (2024c) incorporates in-context learning examples retrieved from a constructed database and CoT prompts that inject domain knowledge of time series. LogPrompt Liu et al. (2024d) explores three prompting strategies for log data: self-prompt, CoT prompt, and in-context prompt, demonstrating that the prompt with CoT techniques outperforms other prompting strategies. The tailored CoT prompt for log data includes a specific task instruction, i.e. “classify the given log entries into normal and abnormal categories”, and step-by-step rules for considering given data as anomalies.
@@ 3.2.1 Detection without LLM Tuning
Anomaly and OOD detection problems can indeed be understood as classification problems. Therefore, pretrained MLLMs like CLIP, with strong zero-shot classification ability, can serve as detectors themselves. By using only ID or normal prompts, CLIP can be leveraged for both OOD and anomaly detection tasks. Despite the promise, existing CLIP-like models perform zero-shot classification in a closed-world setting. That is, it will match an input into a fixed set of categories, even if it is irrelevant Ming et al. (2022). To address this, one approach involves designing effective post-hoc score functions that rely solely on ID or normal class labels. Alternatively, some researchers incorporate anomaly or OOD class information into the text prompts, allowing the model to match OOD or abnormal images to paired prompts.
  • Without Anomaly/OOD Prompts. For anomaly detection, WinCLIP Jeong et al. (2023) initially investigates a one-class design by using only the normal prompt “normal [o]” where [o] represents object-level label, i.e “bottle”, and defining an anomaly score as the similarity between vectors derived from the image encoder and normal prompts. However, this one-class design yields poorer results compared to a simple binary zero-shot framework, CLIP-AC Jeong et al. (2023), which adapts CLIP with two class prompts: “normal [o]” vs. “anomalous [o]”. These results highlight that the one-class design is less effective for anomaly detection, and as a result, anomaly detection research generally does not follow this line with only normal prompts.
    For OOD detection, to address the challenges of using only in-distribution (ID) class information while avoiding the matching of OOD inputs to irrelevant ID classes, one notable approach is the Maximum Concept Matching (MCM) framework proposed by Ming et al. (2022). This method is not limited to CLIP and can be generally applicable to other pre-trained models that promote multi-modal feature alignment. They view the textual embeddings of ID classes as a collection of concept prototypes and define the maximum concept matching (MCM) score based on the cosine similarity between the image feature and the textual feature. Following the idea of MCM, several subsequent works focus on improving OOD detection results by either adding a local MCM score or modifying weights in the original MCM framework, such as Miyai et al. (2023) and Li et al. (2024c).
@@
    For contrasting-based anomaly detection, following the simple binary zero-shot framework, CLIP-AC Jeong et al. (2023), which adapts CLIP with two class prompts: “normal [o]” vs. “anomalous [o]”, many subsequent research emerge. While using the default prompt has demonstrated promising performance, similar to the prompt engineering discussion around GPT-3 Brown et al. (2020), researchers have observed that performance can be significantly improved by customizing the prompt text. Models like WinCLIP Jeong et al. (2023) and AnoCLIP Deng et al. (2023) use a Prompt Ensemble technique to generate all combinations of pre-defined lists of state words per label and text templates. After generating all combinations of states and templates, they compute the average of text embeddings per label to represent the normal and anomalous classes. In practice, more descriptions in prompts do not always yield better performance. Therefore, CLIP-AD Chen et al. (2023b) proposes Representative Vector Selection (RVS), from a distributional perspective for the design of the text prompt, broadening research opportunities beyond merely crafting adjectives.
@@ 3.2.2 Detection with LLM Tuning
  • LLM Adapter-Tuning. Adapter-tuning methods involve integrating additional components or layers into the model architecture to facilitate better alignment or localization Hu et al. (2023). This approach is significantly useful for anomaly detection task, because CLIP was originally designed for classifying the semantics of objects in the scene, which does not align well with the sensory anomaly detection task where both normal and abnormal samples are often from the same class of object. To reconcile this, InCTRL Zhu and Pang (2024) includes a tunable adapter layer to further adapt the image representations for anomaly detection. To better adapt to medical image anomaly detection, MVFA Huang et al. (2024a) proposes a multi-level visual feature adaptation architecture to align CLIP’s features with the requirements of anomaly detection in medical contexts. This is achieved by integrating multiple residual adapters into the pre-trained visual encoder, guided by multi-level, pixel-wise visual-language feature alignment loss functions.
@@
    Instead of focusing on leveraging OOD information into the text encoder, some methods aim to perform prompt tuning to optimize word embeddings for ID labels and then use the MCM score as the detection criterion. MCM-PEFT Ming and Li (2024) demonstrates that simply applying prompt tuning for CLIP on few-shot ID datasets can significantly improve detection accuracy. However, a primary limitation of this approach is its exclusive reliance on the features of ID classes, leading to incorrect detection when input images share a high visual similarity with the class in the prompt. To address this, LoCoOp Miyai et al. (2024c) treats such ID-irrelevant nuisances as OOD and learns to push them away from the ID class text embeddings, preventing the model from producing high ID confidence scores for the OOD features. Additionally, Lafon et al. (2024) enhances detection capabilities by learning a diverse set of prompts utilizing both global and local visual representations. For anomaly detection which emphasizes more on learning local features, AnomalyCLIP Zhou et al. (2024) aims to learn object-agnostic text prompts that capture generic normality and abnormality in images, allowing the model to focus on abnormal regions rather than object semantics.
4 LLMs for Generation
In this section, we review methods that leverage LLMs as generative tools for enhancing anomaly and OOD detection. LLMs use their extensive pre-trained knowledge to generate augmented data, such as embeddings, pseudo labels, and textual descriptions, which improve detection performance. Furthermore, due to their ability to understand and generate human-like text, LLMs have been explored for providing insightful explanations and analyses of detection results, aiding in interpretation, planning, and decision-making. These methods are classified into two main approaches: Augmentation-centric Generation and Explanation-centric Generation.
4.1 Augmentation-centric Generation
LLMs serve as effective tools for data augmentation in anomaly and OOD detection tasks by generating textual embeddings, pseudo labels, and descriptive text. The extensive pre-training of LLMs on large datasets and the autoregressive training objective endow them with superior generative capabilities. These capabilities allow LLMs to produce high-quality embeddings, create synthetic labels, and provide additional descriptive information, ultimately boosting the performance and robustness of detection models.
4.1.1 Text Embedding-based Augmentation
LLMs are highly effective feature extractors, producing meaningful embeddings that can be used in detection tasks. This augmentation enables detection models to capture more subtle patterns and distinctions, leading to more accurate and robust detection performance. For instance, Hadadi et al. (2024) and Qi et al. (2023) fine-tune pre-trained GPT models on log data and use the extracted semantic embeddings as key components for future anomaly detection.
For OOD detection in text data, a common approach involves using encoder-only LLMs to generate sentence representations that are used to compute OOD confidence scores Liu et al. (2024a). These models are typically fine-tuned on ID data, and OOD detectors are applied to the generated representations. Recently, there has been a shift toward leveraging larger language models with decoder architectures, which provide enhanced capabilities in refining textual representations. Liu et al. (2024a) explore the use of decoder-only LLMs, such as LLaMa, incorporating fine-tuning techniques like LoRA to reduce the additional parameters. Their findings demonstrate that fine-tuned LLMs, combined with customized OOD scoring functions, significantly improve OOD detection performance. A key advantage of decoder-based LLMs is their autoregressive ability, which allows for more effective handling of sequential data. Building on this, Zhang et al. (2024a) propose using the likelihood ratio between a pre-trained and fine-tuned LLM as a criterion for OOD detection, leveraging the deep contextual understanding embedded within LLMs for text data.
4.1.2 Pseudo Label-based Augmentation
The emergent capabilities of LLMs provide a promising approach for generating high-quality synthetic datasets, including pseudo labels for OOD samples. A significant challenge in OOD detection is the lack of labeled OOD data, which can limit model performance. Traditionally, obtaining OOD labels requires extensive human effort, but LLMs can mitigate this by generating pseudo-OOD labels through carefully designed prompts. These pseudo labels augment existing ID data and enhance the distinction between ID and OOD samples.
For example, EOE Cao et al. (2024) and PCC Huang et al. (2024b) prompt LLMs to generate visually similar OOD class labels, which are then used to define a new scoring function. This approach significantly outperforms methods relying solely on known ID labels. TOE Park et al. (2023) further evaluates the generation of pseudo-OOD labels at three verbosity levels—word-level, description-level, and caption-level—using BERT, GPT-3, and BLIP-2, respectively. Results indicate that caption-level pseudo-OOD labels generated by BLIP-2, which incorporates both semantic and visual understanding, perform the best. In text data, CoNAL Xu et al. (2023) prompts LLMs to extend closed-set labels with novel examples and generates a comprehensive set of probable OOD samples. By applying contrastive confidence loss during training, the model achieves high accuracy on the ID training set while maintaining lower confidence on the generated OOD examples.
4.1.3 Textual Description-based Augmentation
In addition to generating pseudo labels, LLMs are also used to generate textual descriptions of both known ID classes and potential OOD samples. For example, TagFog Chen et al. (2024) employs a Jigsaw strategy to generate fake OOD samples and prompts ChatGPT to create detailed descriptions for each ID class, guiding the training of the image encoder in CLIP for OOD detection. In anomaly detection tasks, it is essential for LLMs to recognize the close correlation between normal images and their respective prompts, while identifying a more distant association with abnormal prompts. This requires detailed and nuanced descriptions of normal and anomalous stages of objects. ALFA Zhu et al. (2024) formulates prompts for LLMs to describe both normal and abnormal features for each class, and these descriptions are then used to improve the detection of abnormal objects. To avoid LLM hallucination, Dai et al. (2023) introduce a consistency-based uncertainty calibration method, where LLMs describe visual features for distinguishing categories in images, and the confidence score of each generation is estimated accordingly.
Figure 3: The illustration of four approaches in (§4): (a) Text Embedding-based Augmentation; (b) Pseudo Label-based Augmentation; (c) Textual Description-based Augmentation; and (d) Explanation-centric Generation.
4.2 Explanation-centric Generation
Beyond augmentation, LLMs’ powerful reasoning and natural language generation abilities allow them to provide insightful explanations for anomaly and OOD detection outcomes. These explanations are especially important in safety-critical domains, such as autonomous driving, where transparency and interpretability are crucial.